answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of spinning an odd number 8 1 7 2 6 5 4 3?

8 numbers in total, 4 of those numbers are odd (1, 7, 5 and 3)

so 4/8

= 1/2

aka 50%, 0.5

How many permutations are in the word math?

Since the word MATH does not have any duplicated letters, the number of permutations of those letters is simply the number of permutations of 4 things taken 4 at a time, or 4 factorial, or 24.

What is the variance of the uniform distribution function?

the variance of the uniform distribution function is 1/12(square of(b-a)) and the mean is 1/2(a+b).

How many marbles fit into an empty paper bag?

It depends on the size of the marbles and on the size of the paper bag, neither of which was given in the question. Please restate the question.

How do you work out the probability of dice problems?

Count the number of permutations of the expected results and divide by the number of permutations of the possible results. This is standard probability theory, and it applies to everything in probability, not just dice.

For instance, with two dice, there are 36 possible permutations, while there is only 1 permutation that adds up to a sum of 2, though there are 6 permutations that add up to a sum of 7. As a result, the probability of rolling a sum of 2 is 1 in 36, while rolling a 7 is 6 in 36, or 1 in 6.

What are the odds of your house catching on fire?

Naturally if left undisturbed not all that high. But this is one of those things that changes wildly based on other variables. Are there kids in the house, do you live in a "bad" part of town, do you live in a country experiencing war or riots, is the wiring in the house up to code, is the house in an open field, is the house in an area known for wildfires, how much rain has there been in recent times, are any members of the house convicted arsonists, ok by now if you're still reading this you should get the idea.

That being said, if this is secretly an insurance question. Yes you should get homeowners coverage that covers fire.

Is the standard deviation dependent on the mean?

Yes. Consider the definition of the standard deviation. It is the square root of the variance from the mean. As a result, it can be said that the standard deviation is dependent on the mean.

If you were to flip a coin 4 times what is the probability of getting heads at least one time?

As the question is "what is the probability of getting at least one head" the correct way to answer this is to ask what is the probability of not getting any heads and then subtract this from 1.

The probability of not getting a head in 4 flips = 0.54 (i.e. 0.5 * 0.5 * 0.5 * 0.5) = 1/16.

Therefore the probability of getting at least one head is 1 - 1/16 = 15/16.

What is the addition rule of probability?

The addition rule of probability states that the probability that one or the other will happen is the probability of one plus the probability of the other. This rule only applies to mutually exclusive events. For example, the probability that a dice roll will be a 3 is 1/6. The probability that the dice roll will be even is 1/2. These are mutually exclusive events as the dice cannot be both 3 and even. Thus the probability of the dice roll coming up either a 3, or even, is 1/2 + 1/6 = 2/3.

What is the probability of drawing a red or a 5 ball out of a bag that has 4 red balls numbered 1 through 4 and 3 blue balls numbered 3 through 5?

There are 7 balls in total.

4 are red and 1 is the blue "5" ball. Thus there are five balls out of the seven which meet the criteria and therefore the odds must be 5/7.

The probabilities are mutually exclusive (i.e. you can not both draw a red ball and the "5" ball at the same time as the "5" is blue) so you just add the probabilities of each together.

Probability of drawing a red ball = 4/7 (4 of the 7 balls are red)

Probability of drawing the "5" ball = 1/7 (1 of the 7 balls is the "5" ball)

Overall probability = 4/7 + 1/7 = 5/7

How do you calculate standard deviation with the help of z-score?

A z-score cannot help calculate standard deviation. In fact the very point of z-scores is to remove any contribution from the mean or standard deviation.

What is the probability of not rolling a sum of 7 three times in a row?

Possible outcomes for one cube = 6 .

Possible outcomes for 2 cubes = 6 x 6 = 36 .

Ways to roll a sum of 7 = 6 ways,

1, 6 . . . 2, 5 . . . 3, 4

6, 1 . . . 5, 2 . . . 4, 3

Probability of rolling a 7 on the first roll = 6/36 = 1/6 .

Probability of rolling a 7 on the second roll = 6/36 = 1/6 .

Probability of rolling a 7 on the third roll = 6/36 = 1/6 .

Probability of rolling a sum of 7 three times in a row = (1/6 x 1/6 x 1/6) = (1/6)3 = 1/216 = 0.00463 = 0.463%

Probability of not rolling a sum of 7 three times in a row = 1 - (probability of doing it) = 99.537%

How do you do the chapter 4 frappy?

read chapter 4 in your textbook, take notes, practice some problems an then try to do the frappy

What is the probability of picking an ace from a pack of playing cards?

The probability of drawing an Ace from a standard deck of 52 cards is 4 in 52, or 1 in 13, or about 0.07692.

What is the probability of getting court card?

The probability of drawing a court card (a jack, queen, or king) from a standard deck of 52 cards is 12 in 52, or 3 in 13, or about 0.2308.

If you count the ace as a court card, then the probability is 16 in 52, or 4 in 13, or about 0.3077.

A coin is tossed 4 times. What is the probability of getting tails 4 times in a row?

The probability to get tails once is 1/2 (for a fair coin)

The probability to get tails twice = the probability to get it once x the probability to get it a second time

The probability to get tails 4 times in a row is (1/2)4=1/16

The probability to get tails n times in a row is (1/2)n=1/2n

The same thing is also true for heads (same probability: 1/2 each time)

When a single card is drawn from an ordinary 52-card deck find the probability of getting a diamond?

There are 13 diamonds in a 52-card deck

so the probability to get a diamond is 13/52=1/4=0.25

What is the probability of drawing a black card from a 52 card deck and flipping heads on a coin?

Half the cards in a standard pack are black.

Therefore the probability of drawing a black card is 1/2.

Half the sides of a coin are "heads" so again the probability is 1/2.

Therefore the probability you will both draw a black card and flip heads = 1/2 * 1/2 = 1/4.

What is the probability of being dealt 3 cards and all of them being diamonds?

There are 13 diamonds. Three cards are dealt. The probability of all of them being diamond is (13/52)(12/51)(11/50) = 1716/132600 = 11/850

What is symmetrical normal distribution?

The Normal distribution is, by definition, symmetric. There is no other kind of Normal distribution, so the adjective is not used.

When rolling two dice what are the odds of rolling both even numbers?

The probability of rolling two even numbers on two standard dice is 0.52, or 0.25.

What is this answer a coin is tossed and a number cube is rolled Find P heads then greater than 4?

With one toss of a coin, there can be at most 1 head. So the probability of 4 or more heads is very definitely 0.

When a single card is drawn drawn from an ordinary 52 card deck find the probability of getting either a king or a queen?

There are four kings and four queens in a standard 52 card deck. The probability, then, of drawing a king or a queen is 8 in 52, or 2 in 13, or about 0.1538.

Trending Questions
What are the chances of rolling double 6 3 times in a row? Does a multiple choice test question asking for the MOST likely answer eliminate all of the above as a correct answer? What word tells how many standard deviations a measurement is away from the mean? What are the applications of binomial distribution? If two dice were rolled what is the probability of a sum of 5 being rolled? How many ways can the letters of the word hospital be arranged? What is the probability of getting an ace of hearts and a king of hearts from a regular deck of cards without jokers? What is the probability that the arrow will land on either Saturday or Sunday both times? How many bit strings of length ten both begin and end with 1? If a bag contains 20 chips labeled with the numbers 1 through 20 and one chip is selected at random. what is the probability that the chip has the number 17 on it? Why normal distribution is better than other distributions? What chart shows the probability of surviving to any given age? What is a. p. b? What is the difference between a sample and sampling? How many seven's are jacks in a deck of cards? What probability of students? What is the probability that a white sock would be pulled from a drawer that has all white socks in it? How likely are you to pull a red or yellow marble out of a bag of 3 green 8 red 8 yellow and 3 black marbles? If two events cannot occur at the same time those events are considered to be? What is the probability of choosing a blue or green shirt if there's 6?